If h(x) = f(x) + 3, what would the new coordinates of P be after the shift? Give answer in (x, y) form.
P=(2,4)
ii dont understand what its asking? can someone please explain

Answer by vleith(2517)

You are given two functions, f(x) and h(x).
You are given one point that satisfies the function f(x). That point is (2,4).
You are told that h(x) = f(x) + 3
When you think about what a function does, you need to understand that a function 'maps' one value to another. Both f(x) and h(x) map values of x to values of y.
For f(x), you are not told what the function is. You are only told that the point P satisfies that function. What that means is the function f(x), when given x=2, returns y=4.
f(2) = 4
So you get point P (x,y) = (2,4).
With all that background, you are asked to find a point on h(x).
h(x) = f(x) + 3
When x = 2, f(x) = 4, so
h(2) = f(2) + 3
f(2) = 4 (remember point P above?)
h(2) = 4 + 3
h(2) = 7
So the new point is (2,7)